# Repeated Padding for Sequential Recommendation

### v2

@@ 3.5. Implementation Details
Dataset Metric Original Ran SW Ran-S CMR CMRSI RepPad Dataset Metric Original Ran SW Ran-S CMR CMRSI RepPad
Toys Hit@10 0.0256 0.0306 0.0342 0.0363 0.0361 0.0411 0.0434 Toys Hit@10 0.0716 0.0600 0.0522 0.0660 0.0634 0.0713 0.0843
Hit@20 0.0432 0.0491 0.0557 0.0594 0.0585 0.0629 0.0652 Hit@20 0.0965 0.0875 0.0807 0.0950 0.0952 0.1039 0.1142
NDCG@10 0.0121 0.0158 0.0169 0.0192 0.0181 0.0203 0.0229 NDCG@10 0.0396 0.0321 0.0292 0.0366 0.0347 0.0400 0.0476
NDCG@20 0.0166 0.0205 0.0223 0.0248 0.0238 0.0267 0.0284 NDCG@20 0.0471 0.0390 0.0363 0.0439 0.0428 0.0482 0.0551
Beauty Hit@10 0.0327 0.0370 0.0365 0.0419 0.0398 0.0463 0.0468 Beauty Hit@10 0.0637 0.0553 0.0542 0.0563 0.0562 0.0603 0.0726
Hit@20 0.0551 0.0581 0.0604 0.0650 0.0673 0.0712 0.0732 Hit@20 0.0942 0.0850 0.0846 0.0882 0.0876 0.0940 0.1034
NDCG@10 0.0162 0.0191 0.0176 0.0203 0.0196 0.0224 0.0234 NDCG@10 0.0331 0.0285 0.0270 0.0289 0.0290 0.0316 0.0390
NDCG@20 0.0219 0.0241 0.0236 0.0255 0.0266 0.0290 0.0300 NDCG@20 0.0408 0.0359 0.0346 0.0369 0.0639 0.0401 0.0468
Sports Hit@10 0.0143 0.0163 0.0165 0.0194 0.0186 0.0176 0.0204 Sports Hit@10 0.0299 0.0330 0.0315 0.0308 0.0332 0.0318 0.0369
Hit@20 0.0249 0.0267 0.0301 0.0319 0.0309 0.0296 0.0338 Hit@20 0.0447 0.0520 0.0496 0.0514 0.0530 0.0504 0.0553
NDCG@10 0.0065 0.0080 0.0078 0.0096 0.0093 0.0088 0.0100 NDCG@10 0.0160 0.0177 0.0165 0.0154 0.0173 0.0173 0.0196
NDCG@20 0.0091 0.0106 0.0112 0.0127 0.0124 0.0117 0.0134 NDCG@20 0.0197 0.0225 0.0211 0.0206 0.0223 0.0212 0.0242
Yelp Hit@10 0.0172 0.0184 0.0195 0.0210 0.0203 0.0214 0.0223 Yelp Hit@10 0.0282 0.0316 0.0279 0.0364 0.0278 0.0310 0.0359
Hit@20 0.0314 0.0347 0.0349 0.0380 0.0375 0.0389 0.0407 Hit@20 0.0443 0.0516 0.0483 0.0620 0.0480 0.0523 0.0587
NDCG@10 0.0083 0.0085 0.0090 0.0096 0.0103 0.0101 0.0105 NDCG@10 0.0142 0.0162 0.0137 0.0184 0.0136 0.0156 0.0177
NDCG@20 0.0118 0.0126 0.0133 0.0142 0.0139 0.0147 0.0151 NDCG@20 0.0183 0.0212 0.0189 0.0248 0.0187 0.0210 0.0234
Home Hit@10 0.0057 0.0064 0.0072 0.0087 0.0078 0.0094 0.0108 Home H@10 0.0155 0.0156 0.0166 0.0208 0.0167 0.0173 0.0204
H@20 0.0106 0.0120 0.0141 0.0151 0.0149 0.0159 0.0182 H@20 0.0244 0.0247 0.0270 0.0336 0.0275 0.0276 0.0311
N@10 0.0027 0.0030 0.0034 0.0042 0.0037 0.0044 0.0054 N@10 0.0083 0.0086 0.0089 0.0110 0.0092 0.0099 0.0109
N@20 0.0039 0.0045 0.0051 0.0059 0.0054 0.0060 0.0073 N@20 0.0101 0.0105 0.0111 0.0141 0.0113 0.0122 0.0137
3.6. Overall Improvements
The experimental results of original sequential recommendation models and adding our RepPad are presented in Table 4 and Table 5. These tables show that our method can significantly improve the performance of recommendation models on short-sequence datasets, but our method may become ineffective for long-sequence datasets. These results validate our conjecture in Section 2.5.2.
(1) As a whole, the RNN- and CNN-based models perform comparably. NARM is usually the best performer, which verifies the effectiveness of combining attention mechanisms and RNN architectures. The Transformer-based models perform better overall, especially SASRec and STOSA. This shows the ability of the Transformer architecture to model user sequence patterns. FMLP-Rec performs slightly less well, which we believe may be related to the fact that the pure MLP architecture cannot accurately capture the fine-grained relationship between each user interaction and historical behavior as the Transformer does.
(2) As demonstrated in Table 4, after adding our RepPad to the original model, the performance of all models is significantly improved. The average recommendation performance improvement is up to 60.3% on GRU4Rec, 24.3% on SASRec, and 22.8% on STOSA. In particular, Our method gives GRU4Rec a boost of up to 110.00% Hit@5 and 122.22% for NDCG@5 on the Home dataset. The above experimental results validate the effectiveness of our method on short-sequence datasets. It also shows that RepPad can be applied to various types of sequential models. Considering that in real-world scenarios, the vast majority of user history records are short sequences (Liu et al., 2021b; Jiang et al., 2021), our approach has great potential for application.
(3) From Table 5, we can observe that RepPad can only achieve less improvement on some models for the long-sequence dataset and even cause performance degradation on some methods. These results verify our conjecture in Section 2.5.2. Even though we have set the maximum sequence length to 200 for long-sequence datasets, most user sequences still do not have enough space left for repeated padding. At the same time, repeated padding may interfere with the model’s ability to capture long-term user preferences and impair recommendation performance. We also note that the Transformer-based model still outperforms other types of models overall, further demonstrating the power of its architecture.
Table 7. Performance comparison between RepPad and training-required data augmentation methods. The best performance of is bolded and the second best is underlined. The backbone network is SASRec.
Dataset Metric Original ASReP DiffuASR CL4SRec RepPad
Toys Hit@10 0.0716 0.0755 0.0783 0.0804 0.0843
Hit@20 0.0965 0.1021 0.1085 0.1052 0.1142
NDCG@10 0.0396 0.0423 0.0453 0.0447 0.0476
NDCG@20 0.0471 0.0506 0.0535 0.0520 0.0551
Beauty Hit@10 0.0637 0.0664 0.0679 0.0686 0.0726
Hit@20 0.0942 0.0959 0.0966 0.0982 0.1034
NDCG@10 0.0331 0.0351 0.0372 0.0366 0.0390
NDCG@20 0.0408 0.0422 0.0454 0.0449 0.0468
Sports Hit@10 0.0299 0.0324 0.0375 0.0339 0.0369
Hit@20 0.0447 0.0503 0.0566 0.0521 0.0553
NDCG@10 0.0160 0.0166 0.0169 0.0172 0.0196
NDCG@20 0.0197 0.0201 0.0214 0.0225 0.0231
Yelp Hit@10 0.0282 0.0323 0.0345 0.0385 0.0359
Hit@20 0.0443 0.0539 0.0566 0.0621 0.0587
NDCG@10 0.0142 0.0153 0.0169 0.0182 0.0177
NDCG@20 0.0183 0.0207 0.0228 0.0251 0.0234
Home Hit@10 0.0155 0.0176 0.0198 0.0220 0.0204
Hit@20 0.0244 0.0285 0.0301 0.0329 0.0311
NDCG@10 0.0083 0.0098 0.0113 0.0122 0.0116
NDCG@20 0.0101 0.0121 0.0144 0.0149 0.0137
Figure 2. The result of the ablation study.
@@ 3.7. Compare with Other Methods
In this subsection, we compare RepPad with different data augmentation methods. For the heuristic methods, we choose the two representative models, SASRec and GRU4Rec, as the backbone network. For methods that require training, we chose the most versatile SASRec since different methods have different restrictions on the type of backbone network.
@@
From Table 6, we observe that RepPad outperforms existing heuristic data augmentation methods in most cases. CMRSI performs the best among the baselines, thanks to the five well-designed data augmentation operators that help discover more preference information while mitigating data sparsity. Ran-S also achieves competitive performance in some examples due to its augmentation using items in the original sequence. We also observe that in some cases, the model performance degraded after using data augmentation methods, which may be related to the fact that these methods introduce too much noise that destroys the integrity of the original sequence. RepPad utilizes the free data space and uses the original sequences for padding, which improves model performance while retaining the user’s complete preference knowledge. From Table 7, RepPad still achieves satisfactory results in the face of data augmentation methods that require training. Among the baseline methods, DiffuASR and CL4SRec leverage the diffusion model and sequence-level augmentation with contrastive learning to achieve competitive performance in some cases. Surprisingly, RepPad achieves the best or second-best performance in all cases without including any parameters and training process.
@@
Overall, RepPad proved its superiority by achieving highly competitive results without including any parameters or training processes. Also, our approach is not limited by the backbone network and can be seamlessly inserted into most existing sequential recommendation models.

### v3

@@ 4.2. Overall Improvements
Dataset Metric Original Ran SW Ran-S CMR CMRSI RepPad RepPad+ Dataset Metric Original Ran SW Ran-S CMR CMRSI RepPad RepPad+
Toys H@10 0.0256 0.0306 0.0342 0.0363 0.0361 0.0411 0.0434 0.0482 Toys H@10 0.0716 0.0600 0.0522 0.0660 0.0634 0.0713 0.0843 0.0866
H@20 0.0432 0.0491 0.0557 0.0594 0.0585 0.0629 0.0652 0.0716 H@20 0.0965 0.0875 0.0807 0.0950 0.0952 0.1039 0.1142 0.1172
N@10 0.0121 0.0158 0.0169 0.0192 0.0181 0.0203 0.0229 0.0269 N@10 0.0396 0.0321 0.0292 0.0366 0.0347 0.0400 0.0476 0.0512
N@20 0.0166 0.0205 0.0223 0.0248 0.0238 0.0267 0.0284 0.0328 N@20 0.0471 0.0390 0.0363 0.0439 0.0428 0.0482 0.0551 0.0589
Beauty H@10 0.0327 0.0370 0.0365 0.0419 0.0398 0.0463 0.0468 0.0505 Beauty H@10 0.0637 0.0553 0.0542 0.0563 0.0562 0.0603 0.0726 0.0762
H@20 0.0551 0.0581 0.0604 0.0650 0.0673 0.0712 0.0732 0.0756 H@20 0.0942 0.0850 0.0846 0.0882 0.0876 0.0940 0.1034 0.1089
N@10 0.0162 0.0191 0.0176 0.0203 0.0196 0.0224 0.0234 0.0263 N@10 0.0331 0.0285 0.0270 0.0289 0.0290 0.0316 0.0390 0.0424
N@20 0.0219 0.0241 0.0236 0.0255 0.0266 0.0290 0.0300 0.0326 N@20 0.0408 0.0359 0.0346 0.0369 0.0639 0.0401 0.0468 0.0507
Sports H@10 0.0143 0.0163 0.0165 0.0194 0.0186 0.0176 0.0204 0.0232 Sports H@10 0.0299 0.0330 0.0315 0.0308 0.0332 0.0318 0.0369 0.0434
H@20 0.0249 0.0267 0.0301 0.0319 0.0309 0.0296 0.0338 0.0383 H@20 0.0447 0.0520 0.0496 0.0514 0.0530 0.0504 0.0553 0.0641
N@10 0.0065 0.0080 0.0078 0.0096 0.0093 0.0088 0.0100 0.0115 N@10 0.0160 0.0177 0.0165 0.0154 0.0173 0.0173 0.0196 0.0239
N@20 0.0091 0.0106 0.0112 0.0127 0.0124 0.0117 0.0134 0.0153 N@20 0.0197 0.0225 0.0211 0.0206 0.0223 0.0212 0.0242 0.0291
Yelp H@10 0.0172 0.0184 0.0195 0.0210 0.0203 0.0214 0.0223 0.0274 Yelp H@10 0.0282 0.0316 0.0279 0.0364 0.0278 0.0310 0.0359 0.0374
H@20 0.0314 0.0347 0.0349 0.0380 0.0375 0.0389 0.0407 0.0474 H@20 0.0443 0.0516 0.0483 0.0620 0.0480 0.0523 0.0587 0.0601
N@10 0.0083 0.0085 0.0090 0.0096 0.0103 0.0101 0.0105 0.0129 N@10 0.0142 0.0162 0.0137 0.0184 0.0136 0.0156 0.0177 0.0190
N@20 0.0118 0.0126 0.0133 0.0142 0.0139 0.0147 0.0151 0.0179 N@20 0.0183 0.0212 0.0189 0.0248 0.0187 0.0210 0.0234 0.0255
Home H@10 0.0057 0.0064 0.0072 0.0087 0.0078 0.0094 0.0108 0.0121 Home H@10 0.0155 0.0156 0.0166 0.0208 0.0167 0.0173 0.0204 0.0211
H@20 0.0106 0.0120 0.0141 0.0151 0.0149 0.0159 0.0182 0.0201 H@20 0.0244 0.0247 0.0270 0.0336 0.0275 0.0276 0.0311 0.0324
N@10 0.0027 0.0030 0.0034 0.0042 0.0037 0.0044 0.0054 0.0063 N@10 0.0083 0.0086 0.0089 0.0109 0.0092 0.0099 0.0110 0.0124
N@20 0.0039 0.0045 0.0051 0.0059 0.0054 0.0060 0.0073 0.0083 N@20 0.0101 0.0105 0.0111 0.0141 0.0113 0.0122 0.0137 0.0145
@@ 4.3. Compare with Other Methods
In this subsection, we compare RepPad+ with different data augmentation methods. For the heuristic methods, we choose the two representative models, SASRec and GRU4Rec, as the backbone network. For methods that require training, we chose the most versatile SASRec since different methods have different restrictions on the type of backbone network. For example, the ASReP is based on Transformer architecture and can not be applied to GRU4Rec. We present the results in Table 6 and 7.
Table 7. Performance comparison between previous work RepPad, our method RepPad+, and training-required data augmentation methods. The best performance is bolded and the second best is underlined. The backbone network is SASRec.
Dataset Metric Original ASReP DiffuASR CL4SRec RepPad RepPad+
Toys H@10 0.0716 0.0755 0.0783 0.0804 0.0843 0.0866
H@20 0.0965 0.1021 0.1085 0.1052 0.1142 0.1172
N@10 0.0396 0.0423 0.0453 0.0447 0.0476 0.0512
N@20 0.0471 0.0506 0.0535 0.0520 0.0551 0.0589
Beauty H@10 0.0637 0.0664 0.0679 0.0686 0.0726 0.0762
H@20 0.0942 0.0959 0.0966 0.0982 0.1034 0.1089
N@10 0.0331 0.0351 0.0372 0.0366 0.0390 0.0424
N@20 0.0408 0.0422 0.0454 0.0449 0.0468 0.0507
Sports H@10 0.0299 0.0324 0.0375 0.0339 0.0369 0.0434
H@20 0.0447 0.0503 0.0566 0.0521 0.0553 0.0641
N@10 0.0160 0.0166 0.0169 0.0172 0.0196 0.0239
N@20 0.0197 0.0201 0.0214 0.0225 0.0242 0.0291
Yelp H@10 0.0282 0.0323 0.0345 0.0385 0.0359 0.0374
H@20 0.0443 0.0539 0.0566 0.0621 0.0587 0.0601
N@10 0.0142 0.0153 0.0169 0.0182 0.0177 0.0190
N@20 0.0183 0.0207 0.0228 0.0251 0.0234 0.0255
Home H@10 0.0155 0.0176 0.0198 0.0220 0.0204 0.0211
H@20 0.0244 0.0285 0.0301 0.0329 0.0311 0.0324
N@10 0.0083 0.0098 0.0113 0.0122 0.0116 0.0124
N@20 0.0101 0.0121 0.0124 0.0149 0.0137 0.0145
@@
From Table 6, we observe that RepPad+ outperforms existing heuristic data augmentation methods and the previous work RepPad. Among the baseline methods, CMRSI usually performs the best, thanks to the five well-designed data augmentation operators that help mine more preference information while mitigating data sparsity. Ran-S also achieves competitive performance in some examples due to its augmentation using items in the original sequence. In some cases, the model performance degraded after using heuristic data augmentation methods, which may be related to the fact that these methods introduce too much noise that destroys the integrity of the original sequence. In most cases, RepPad can achieve the second-best performance, while RepPad+ can achieve the best. From Table 7, RepPad+ still achieves satisfactory results in the face of data augmentation methods that require training. Among the baseline methods, DiffuASR and CL4SRec leverage the diffusion model and sequence-level augmentation with contrastive learning to achieve competitive performance in some cases. Surprisingly, RepPad+ achieves the best or second-best performance in all cases without including any parameters and training process.
@@
Overall, RepPad+ proved its superiority by achieving highly competitive results without including any parameters or training processes. Compared to RepPad, RepPad+ also demonstrated the importance of augmenting the medium-length sequences. Also, our approach is not limited by the backbone network and can be seamlessly inserted into most existing sequential recommendation models.
4.4. Improvement on Long and Short Sequences
Following BiCAT (Jiang et al., 2021), we explore the effectiveness of our methods on short and long sequences in the same dataset and report the results in Table 8 and 9. A long sequence is a sequence whose length $\left|s_{u}\right|$ fulfill the condition $20<\left|s_{u}\right|$. Since we perform 5-core filtering during the dataset, a short sequence is an interaction sequence of length 5 (i.e., $5=\left|s_{u}\right|$). From the tables, we observe a similar trend as in Table 5. With the addition of the RepPad, model performance can often achieve significant gains on short sequences while achieving limited performance gains on long sequences, and in many cases, even producing sizable performance degradation. Compared to RepPad, RepPad+ improves the model’s performance on long sequences by augmenting medium-length sequences. For items in long sequences, the model can learn their representations more accurately, and these items are likely to appear in short sequences, so RepPad+ improves the model’s recommendation performance for short sequences as well. Overall, RepPad+ makes up for RepPad’s poor performance on long sequences and further improves the overall recommendation accuracy of the model. Comparing the results of our method on long and short sequences, our method is highly applicable to mitigating the cold start problem, as well as facilitating the learning of long-tailed item representations.
Table 8. Model performance on different sequences. The ‘w/ RP’ and ‘w/ RP+’ represent adding RepPad and our RepPad+ to the corresponding model, respectively. The ‘Imp-O’ and ‘Imp-R’ represent the relative improvements of RepPad+ over the original model and the RepPad, respectively. The backbone network is GRU4Rec.
Sequence Length Long ($20<\left|s_{u}\right|\leq 50$) Short ($\left|s_{u}\right|=5$)
Dataset Metric Original w/ RP Imp-O w/ RP+ Imp-O Imp-R Original w/ RP Imp-O w/ RP+ Imp-O Imp-R
Toys H@10 0.0314 0.0392 24.84% 0.0597 90.13% 52.30% 0.0251 0.0354 41.04% 0.0419 66.93% 18.36%
H@20 0.0769 0.0689 -10.40% 0.1052 36.80% 52.69% 0.0426 0.0534 25.35% 0.0642 50.70% 20.22%
N@10 0.0133 0.0207 55.64% 0.0279 109.77% 34.78% 0.0130 0.0199 53.08% 0.0222 70.77% 11.56%
N@20 0.0245 0.0273 11.61% 0.0391 59.85% 43.22% 0.0174 0.0244 40.23% 0.0278 59.77% 13.93%
Beauty H@10 0.0663 0.0889 34.09% 0.1139 71.79% 28.12% 0.0264 0.0350 32.58% 0.0489 85.23% 39.71%
H@20 0.1189 0.1377 15.81% 0.1564 31.54% 13.58% 0.0417 0.0588 41.01% 0.0711 70.50% 20.92%
N@10 0.0383 0.0497 29.77% 0.0587 53.26% 18.11% 0.0124 0.0172 38.71% 0.0293 136.29% 70.35%
N@20 0.0516 0.0618 19.77% 0.0695 34.69% 12.46% 0.0163 0.0231 41.72% 0.0349 114.11% 51.08%
Sports H@10 0.0212 0.0193 -8.96% 0.0198 -6.60% 2.59% 0.0138 0.0203 47.10% 0.0259 87.68% 27.59%
H@20 0.0308 0.0285 -7.47% 0.0353 14.61% 23.86% 0.0243 0.0348 43.21% 0.0414 70.37% 18.97%
N@10 0.0097 0.0088 -9.28% 0.0093 -4.12% 5.68% 0.0063 0.0102 61.90% 0.0128 103.17% 25.49%
N@20 0.0122 0.0114 -6.56% 0.0132 8.20% 15.79% 0.0089 0.0138 55.06% 0.0167 87.64% 21.01%
Yelp H@10 0.0160 0.0176 10.00% 0.0267 66.88% 51.70% 0.0176 0.0235 33.52% 0.0255 44.89% 8.51%
H@20 0.0342 0.0374 9.36% 0.0465 35.96% 24.33% 0.0332 0.0410 23.49% 0.0435 31.02% 6.10%
N@10 0.0081 0.0082 1.23% 0.0115 41.98% 40.24% 0.0085 0.0114 34.12% 0.0125 47.06% 9.65%
N@20 0.0126 0.0131 3.97% 0.0166 31.75% 26.72% 0.0124 0.0161 29.84% 0.0170 37.10% 5.59%
Home H@10 0.0241 0.0213 -11.62% 0.0257 6.64% 20.66% 0.0042 0.0089 111.90% 0.0116 176.19% 30.34%
H@20 0.0408 0.0354 -13.24% 0.0389 -4.66% 9.89% 0.0076 0.0150 97.37% 0.0185 143.42% 23.33%
N@10 0.0115 0.0107 -6.96% 0.0123 6.96% 14.95% 0.0020 0.0048 140.00% 0.0062 210.00% 29.17%
N@20 0.0158 0.0142 -10.13% 0.0157 -0.63% 10.56% 0.0029 0.0064 120.69% 0.0079 172.41% 23.44%
Average - - 6.57% - 34.24% 25.11% - - 55.60% - 93.26% 23.77%
Table 9. Model performance on different sequences. The ‘w/ RP’ and ‘w/ RP+’ represent adding RepPad and our RepPad+ to the corresponding model, respectively. The ‘Imp-O’ and ‘Imp-R’ represent the relative improvements of RepPad+ over the original model and the RepPad, respectively. The backbone network is SASRec.
Sequence Length Long ($20<\left|s_{u}\right|\leq 50$) Short ($\left|s_{u}\right|=5$)
Dataset Metric Original w/ RP Imp-O w/ RP+ Imp-O Imp-R Original w/ RP Imp-O w/ RP+ Imp-O Imp-R
Toys H@10 0.0863 0.0863 0.00% 0.0973 12.75% 12.75% 0.0743 0.0807 8.61% 0.0888 19.52% 10.04%
H@20 0.1334 0.1303 -2.32% 0.1381 3.52% 5.99% 0.0974 0.1075 10.37% 0.1161 19.20% 8.00%
N@10 0.0474 $0$15.61% 0.0546 15.19% 36.50% 0.0435 0.0472 8.51% 0.0543 24.83% 15.04%
N@20 0.0594 0.0511 -13.97% 0.0650 9.43% 27.20% 0.0493 0.0539 9.33% 0.0612 24.14% 13.54%
Beauty H@10 0.1250 0.1261 0.88% 0.1327 6.16% 5.23% 0.0526 0.0623 18.44% 0.0702 33.46% 12.68%
H@20 0.1779 0.1711 -3.82% 0.1902 6.91% 11.16% 0.0797 0.0881 10.54% 0.0987 23.84% 12.03%
N@10 0.0654 0.0693 5.96% 0.0773 18.20% 11.54% 0.0280 0.0346 23.57% 0.0381 36.07% 10.12%
N@20 0.0788 0.0806 2.28% 0.0919 16.62% 14.02% 0.0348 0.0412 18.39% 0.0452 29.89% 9.71%
Sports H@10 0.0408 0.0397 -2.70% 0.0441 8.09% 11.08% 0.0328 0.0396 20.73% 0.0469 42.99% 18.43%
H@20 0.0540 0.0507 -6.11% 0.0606 12.22% 19.53% 0.0510 0.0587 15.10% 0.0706 38.43% 20.27%
N@10 0.0224 0.0223 -0.45% 0.0241 7.59% 8.07% 0.0167 0.0215 28.74% 0.0260 55.69% 20.93%
N@20 0.0258 0.0250 -3.10% 0.0282 9.30% 12.80% 0.0213 0.0263 23.47% 0.0320 50.23% 21.67%
Yelp H@10 0.0271 0.0262 -3.32% 0.0347 28.04% 32.44% 0.0230 0.0318 38.26% 0.0349 51.74% 9.75%
H@20 0.0474 0.0459 -3.16% 0.0545 14.98% 18.74% 0.0347 0.0540 55.62% 0.0599 72.62% 10.93%
N@10 0.0136 0.0137 0.74% 0.0166 22.06% 21.17% 0.0106 0.0154 45.28% 0.0172 62.26% 11.69%
N@20 0.0186 0.0188 1.08% 0.0215 15.59% 14.36% 0.0136 0.0210 54.41% 0.0235 72.79% 11.90%
Home H@10 0.0374 0.0277 -25.94% 0.0344 -8.02% 24.19% 0.0122 0.0195 59.84% 0.0208 70.49% 6.67%
H@20 0.0505 0.0479 -5.15% 0.0521 3.17% 8.77% 0.0188 0.0292 55.32% 0.0323 71.81% 10.62%
N@10 0.0189 0.0137 -27.51% 0.0178 -5.82% 29.93% 0.0065 0.0109 67.69% 0.0117 80.00% 7.34%
N@20 0.0222 0.0209 -5.86% 0.0223 0.45% 6.70% 0.0080 0.0134 67.50% 0.0146 82.50% 8.96%
Average - - -5.40% - 9.82% 16.61% - - 31.99% - 48.12% 12.52%